# Go-or-Grow Models in Biology: a Monster on a Leash

### v1

@@ 7.1   The FKPP case
The analysis of traveling waves for the FKPP equation (1) is a well-researched and well-described textbook method [110, 136, 9, 88, 89]. Consider (1) on the real line and assume the growth function $f$ is of logistic type (2), we introduce the traveling wave coordinate $z=x-ct$ with wave speed $c$, and define a traveling wave solution to be a self-similar solution of the form $m(t,x)=M(z)$. Substituting this ansatz into (1) yields a second-order ODE for $M$ of the form
$-cM^{\prime}=dM^{\prime\prime}+g(M)M,$ (38)
@@ 7.2   The go-or-grow case
Assumption (39) guarantees that ${\bf F}(0,0)={\bf F}(u^{*},v^{*})=(0,0)$ in (10), allowing us to formalize the question of a traveling wave with the following definitions.
@@
Consider the general go-or-grow model (4) and introduce the traveling wave coordinate $z=x-ct$ for a fixed speed $c$. A solution of the form $u(t,x)=U(x-ct)$ and $v(t,x)=V(x-ct)$ is called a traveling wave solution with speed $c$. The traveling wave is called a wavefront if $\big{(}U(z),V(z)\big{)}$ has limits
@@
For the balanced go-or-grow model (7) with $g(n)=1-n$, Pham et al. [121] showed numerically that invasion waves exist with different qualitative behavior. In particular, the wake of the wave might be a homogeneous steady state, a periodic wave train, or even an irregular pattern, as we have shown in Section 5.
More recently, Falcó et al. [40] examined the minimum invasion speed in the case of the total population go-or-grow model (6) with no death rate ($\mu=0$), logistic growth for the proliferating population ($g(n)=1-n$), and $\alpha(n)$ and $\beta(n)$ as nonincreasing and nondecreasing functions, respectively. By assuming $\alpha(n)>0$ for all $n>0$, they ensure the existence of only two equilibria in the system,
@@
    In this case, if $\beta(n)\neq 0$ for $n>0$, the positivity of traveling wave solutions implies that a wave speed faster than $\bar{c}$ is always exhibited.
@@ 7.3   A unifying framework and a general result
$\begin{split}\bullet\,\,&\alpha(u,v)\;\mbox{and }\beta(u,v)\;\mbox{are piecewise continuously differentiable in}\;(u,v)\;\mbox{for}\;{0\leq u\leq u^{*}}\\ &\mbox{and}\;0\leq v\leq v^{*}\mbox{, differentiable at}\;(0,0)\mbox{, and}\;\hskip 213.39566pt\\[2.84544pt] &\hskip 71.13188pt(g(0)-\beta(0,0))(\mu+\alpha(0,0))+\alpha(0,0)\beta(0,0)>0;\end{split}$ (43a)
@@
[…]
We transform system (10) the into wave coordinate $z=x-ct$ and linearize about $(0,0)$ to obtain
@@
[…]
@@
[…]
@@
Although condition (49) looks quite complex, it is already satisfied when ${\bf F}(u,v)$ is twice differentiable in both components [41]. The assumption (48) comes from a variant of the subtangential condition (3), requiring that ${\bf F}(u,v)\leq{\bf F}^{\prime}(0,0)$ along the direction of the eigenvector associated with the principal eigenvalue $\lambda_{A}(\rho)$. With assumptions (39), (43a)–(43b), and (49)–(48), we can show that (4) has traveling wave solutions.
@@
If the general grow-or-grow model (4) satisfies the conditions (39), (43a)–(43b), and (48)–(49), then for each $c<\bar{c}$, it has a nonincreasing wavefront $\big{(}U(x-ct),V(x-ct)\big{)}$ connecting $(u^{*},v^{*})$ to $(0,0)$. However, for any $c\in(0,\bar{c})$, there is no wavefront $\big{(}U(x-ct),V(x-ct)\big{)}$ connecting $(u^{*},v^{*})$ to $(0,0)$.
@@ 7.4   Connection between the spread rate of the FKPP equation and go-or-grow models
Assume $\mu=0$, ${\bf F}({\bf w})$ is given by (11), and $\alpha$, $\beta$, and $g$ satisfy assumptions (2), (3), (39), (43a)–(43b), and (48)–(49). Then the minimum wave speed, $\bar{c}$ of the go-or-grow model (4) given by (47) satisfies the inequality
@@
Since ${\bf F}(u,v)$ satisfies assumptions (39), (43a)–(43b), and (48)–(49), Theorem 8 can be applied, and we have that the minimum wave speed of the corresponding go-or-grow model is given by
@@
Some special cases of this result were already proven in the literature for the constant rates model (8) in [45, 90] and for the total population model (6) in [40]. However, the results provided here are a generalization for the case of the general go-or-grow model (4).
@@ 7.5   Large wave speed scaling
Introducing the traveling wave coordinate, $z=x-ct$, such that $U(z)=u(t,x)$ and $V(z)=v(t,x)$, the general go-or-grow model (4) is written as
@@ 8   Discussion
Focusing on specific analytical properties of go-or-grow models that can be effectively addressed, we extensively reviewed results related to the critical domain problem and the analysis of the existence and minimal speed of traveling wave solutions for the general go-or-grow model and its specialized variants. For the critical domain problem, we revisited classical results concerning the FKPP equation [9, 158] and connected them to recent findings related to the general go-or-grow model [151]. Under specific assumptions about the transition and growth functions, we significantly generalize previous results on the bifurcation between domain size and the existence of non-trivial steady-state solutions. In our examination of traveling wave solutions, we outlined the necessary assumptions that the general go-or-grow model (4) must meet to be classified as a cooperative system. This classification allows us to apply a robust theoretical framework concerning the existence of traveling waves and invasion speed, as established in the literature [41, 163]. Similar to our analysis of the critical domain size, we establish a connection to the FKPP equation, thereby generalizing previous results to the context of the general go-or-grow model (4) introduced here.
@@ Acknowledgements
We are grateful to the Fields Institute in Toronto, who hosted us for a few weeks in Summer and Fall 2024 to work on this project. RT acknowledges support from a Canadian Graduate Scholarship of the Natural Sciences and Engineering Research Council of Canada (NSERC) and Alberta Innovates Graduate Student Scholarship. MC was supported by the National Group of Mathematical Physics (GNFM-INdAM) through the INdAM–GNFM Project (CUP E53C22001930001) “From kinetic to macroscopic models for tumor-immune system competition”, and by the European Union - NextGenerationEU and the MUR-Italian Ministry of Universities and Research through the project PRIN 2022 PNRR Mathematical modeling for a Sustainable Circular Economy in Ecosystems (project code P2022PSMT7, CUP D53D23018960001). TLS acknowledges support from a National Science Foundation (NSF) grant DMS-2151566. TH is supported through a discovery grant of the Natural Science and Engineering Research Council of Canada (NSERC), RGPIN-2023-04269.
@@ Appendix A Appendix: Proofs of Lemma 7 and Theorem 8
The proofs of Lemma 7 and Theorem 8 primarily require showing that the general go-or-grow model (4) with the assumptions (39), (43a)–(43b), and (48)–(49) lives inside the general cooperative theory proposed by [96, 41, 163]. In these works, the authors are concerned the $n$-equation reaction–diffusion system
@@
$\begin{split}\bullet\,\,&{\bf F}\text{ is continuous with }\,{\bf F}(\mathbf{0})={\bf F}({\bf\bar{w}})={\bf 0},\,\text{ and there is no }\,{\bf\nu}\,\text{ such that }{\bf F}({\bf\nu})=\mathbf{0}\,\\ &\text{ and }\,\mathbf{0}<\nu<{\bf\bar{w}};\end{split}$ (58a)
@@
$\begin{split}\bullet\,\,&{\bf F}(\mathbf{w})\,\text{ is piecewisecontinuously differentiable in }\,\mathbf{w}\,\text{ for }\,\mathbf{0}\leq\mathbf{w}\leq{\bf\bar{w}}\,\text{ and differentiable at }\mathbf{0}\,,\\ &\text{ and the matrix }\,{\bf F}^{\prime}(\mathbf{0})\,\text{ is irreducible with a positive eigenvalue;}\end{split}$ (58c)
$\begin{split}\bullet\,\,&\text{There exists }\,a>0\,,\sigma>1\,,\text{and }{\bf r}>0\,\text{ such that }\\[2.84544pt] &\hskip 156.49014pt{\bf F}(\mathbf{w})\geq{\bf F}^{\prime}(\mathbf{0})\mathbf{w}-a\|\mathbf{w}\|^{\sigma}\mathbf{1},\\ &\text{for all }\,\mathbf{0}\leq\mathbf{w}\leq{\bf r},\text{ where }\mathbf{1}\text{ denotes the vector with all entries equal to 1;}\end{split}$ (58d)
$\begin{split}\bullet\,\,&\text{For any }\,\eta>0,\\[2.84544pt] &\hskip 142.26378pt{\bf F}(\min\{\eta\mathbf{w})(\rho),{\bf 1}\}\leq\eta\,D{\bf F}(\mathbf{0})\mathbf{w}(\rho),\,\\ &\text{for all}\,\rho\in(0,\rho^{*}]\,,\text{ where }\,\rho^{*}\,\text{ is thevalue of }\,\rho\,\text{ at which }\,\lambda_{+}(\rho)/\rho\,\text{ attains its infimum.}\end{split}$ (58e)
@@
If the general go-or-grow model (4) satisfies (39), (43a)–(43b), and (48)–(49), then (4) also satisfies the set of assumptions (58).
@@
Condition (58a), and the first part of condition (58c) are directly satisfied by the assumptions (39) and (43a). In addition, during the wavefront calculation (Section 7.3), we computed the eigenvalue of the matrix $A=\rho^{2}D+{\bf F}^{\prime}(0,0)$ (45), which for $\rho=0$ reads
@@
Now, the more technical assumptions (58d) and (58e) need to be satisfied. Considering the first inequality of assumption (49), rearranging terms, and adding $-\mu\,u$ to both sides yields
@@
adding $-\mu\nu_{1}(\rho)$ to both sides, and multiplying by $\eta>0$, we get
@@
Therefore, if (4) satisfies (39), (43a)–(43b), and (48)–(49), then (4) also satisfies (58). ∎
@@
  • (i) $\Phi(\mu)=\dfrac{\lambda_{A}(\mu)}{\mu}$ is decreasing as $\mu$ nears 00 and tends to infinity as $\mu\rightarrow 0$; $\Phi^{\prime}(\mu)$ changes sign at most once in $(0,\infty)$ and $\lim\limits_{\mu\rightarrow\infty}\Phi(\mu)=+\infty$;
@@
  • (iii) For any $c\in(0,\tilde{c})$ and any vector $\mathbf{r}>\mathbf{0}$, there is a positive number $R_{r}$ such that for any $\phi\in\mathcal{C}_{{\bf\bar{w}}}$ with $\phi\leq\mathbf{r}$ on an interval of length $2R_{\mathbf{r}}$, it holds that $\lim\limits_{\begin{subarray}{c}t\rightarrow\infty\\ |x|\leq ct\end{subarray}}{\bf w}(t,x;\phi)={\bf\bar{w}}$;
@@
Applying Lemma 11 and Theorem 12 to the general go-org-grow model (4), with assumptions (39), (43a)–(43b), and (48)–(49), proves Lemma 7 and Theorem 8.

### v2

@@ 7.1   The FKPP case
The analysis of traveling waves for the FKPP equation (1) is a well-researched and well-described textbook method [87, 108, 9, 71, 73]. Consider (1) on the real line and assume the growth function $f$ is of logistic type (2), we introduce the traveling wave coordinate z=xctz=x-ct with wave speed $c$, and define a traveling wave solution to be a self-similar solution of the form $m(t,x)=M(z)$. Substituting this ansatz into (1) yields a second-order ODE for $M$ of the form
$cM^{\prime}=dM^{\prime\prime}+g(M)M,$ (35)
@@ 7.2   The go-or-grow case
[…]
$\hskip 56.9055pt\left\{\begin{array}[]{@{}l@{}}0=-\mu u^{*}-\alpha(u^{*},v^{*})u^{*}+\beta(u^{*},v^{*})v^{*},\\[5.69046pt] \mu u^{*}=g(u^{*}+v^{*})v^{*}.\end{array}\right$. (36e)
Assumptions (36a)-(36b) guarantees that ${\bf F}(0,0)={\bf F}(u^{*},v^{*})=(0,0)$ in (10), allowing us to formalize the question of a traveling wave with the following definitions.
@@
Consider the general go-or-grow model (4) and introduce the traveling wave coordinate z=xctz=x-ct for a fixed speed $c$. A solution of the form u(t,x)=U(xct)u(t,x)=U(x-ct) and v(t,x)=V(xct)v(t,x)=V(x-ct) is called a traveling wave solution with speed $c$. The traveling wave is called a wavefront if $\big(U(z),V(z)\big)$ has limits
@@
For the balanced go-or-grow model (7) with g(n)=1ng(n)=1-n, Pham et al. [97] showed numerically that invasion waves exist with different qualitative behavior. In particular, the wake of the wave might be a homogeneous steady state, a periodic wave train, or even an irregular pattern, as we have shown in Section 5.
More recently, Falcó et al. [34] examined the minimum invasion speed in the case of the total population go-or-grow model (6) with no death rate ($\mu=0$), logistic growth for the proliferating population (g(n)=1ng(n)=1-n), and $\alpha(n)$ and $\beta(n)$ as nonincreasing and nondecreasing functions, respectively. By assuming $\alpha(n)>0$ for all $n>0$, they ensure the existence of only two equilibria in the system,
@@
    In this case, if $\beta(n)\neq 0$ for $n>0$, it has been demonstrated that ensuring the positivity of traveling wave solutions requires a wave speed $c>\bar{c}$. Therefore, for the case where $\beta(0)=0$, the traveling wave solutions exhibit a wave speed greater than $\bar{c}$ meaning that the wave speed is nonlinearly determined.
@@ 7.3   A unifying framework and a general result
[…]
@@
Assumption (40a) gives regularity such that we can take the Jacobian at $(0,0)$ and ensures that the Jacobian at $(0,0)$ has a positive eigenvalue. Note that, if $\mu=0$, the condition $(g(0)-\beta(0,0))(\mu+\alpha(0,0))+\alpha(0,0)\beta(0,0)>0$} is satisfied. The assumption (40b) is sufficient to ensure that (4) is cooperative.
We transform system (10) the into wave coordinate z=xctz=x-ct and linearize about $(0,0)$ to obtain
@@
$\nu(\rho)=\left(\frac{\lambda_{A}(\rho)-g(0)+\beta(0,0)}{\alpha(0,0)},\;1\right)^{T},$ (43)
@@
Let $\alpha$, $\beta$, and $g$ satisfy (36b), (40a), and (40b) and consider the initial condition $\phi=(\phi_{1},\phi_{2})\in\mathcal{C}_{u^{*},v^{*}}=\{(\phi_{1},\phi_{2})\in C(\mathbb{R};\mathbb{R}^{2}):0\leq\phi_{1}(z)\leq u^{*},\;0\leq\phi_{2}(z)\leq v^{*}\;\forall z\in\mathbb{R}\}$. Let $\mathbf{w}(t,x;\phi):=\big(u(t,x;\phi_{1}),v(t,x;\phi_{2})\big)$} be the unique solution of (4) through $\phi$. Then, there exists a real number $c^{*}$ with $0<c^{*}\leq\bar{c}$, such that the following statements are valid:
@@
Although condition (46) looks quite complex, it is already satisfied when ${\bf F}(u,v)$ is twice differentiable in both components [35]. The assumption (45) comes from a variant of the subtangential condition (3), requiring that ${\bf F}(u,v)\leq{\bf F}^{\prime}(0,0)$ along the direction of the eigenvector associated with the principal eigenvalue $\lambda_{A}(\rho)$. With assumptions (36b), (40a)–(40b), and (46)–(45), we can show that (4) has traveling wave solutions.
@@
If the general grow-or-grow model (4) satisfies the conditions (36b), (40a)–(40b), and (45)–(46), then for each $c<\bar{c}$, it has a nonincreasing wavefront $\big(U(x-ct),V(x-ct)\big)$ connecting $(u^{*},v^{*})$ to $(0,0)$. However, for any $c\in(0,\bar{c})$, there is no wavefront $\big(U(x-ct),V(x-ct)\big)$ connecting $(u^{*},v^{*})$ to $(0,0)$.
@@ 7.4   Connection between the spread rate of the FKPP equation and go-or-grow models
Assume $\mu=0$, ${\bf F}({\bf w})$ is given by (11), and $\alpha$, $\beta$, and $g$ satisfy assumptions (2), (3), (36b), (40a)–(40b), and (45)–(46). Then the minimum wave speed, $\bar{c}$ of the go-or-grow model (4) given by (44) satisfies the inequality
@@
Since ${\bf F}(u,v)$ satisfies assumptions (36b), (40a)–(40b), and (45)–(46), Theorem 7 can be applied, and we have that the minimum wave speed of the corresponding go-or-grow model is given by
@@
Some special cases of this result were already proven in the literature for the constant rates model (8) in [37, 74] and for the total population model (6) in [34]. However, the results provided here are a generalization for the case of the general go-or-grow model (4). As an example, in Figure 6 we illustrate the traveling wave profile for the constant rates model (8) (black and red lines) and the FKPP model (1) (blue lines). These numerical simulations show how the FKPP model is at least two times faster than the constant rates go-or-grow model.
Figure 6: Traveling wave profiles for models (1) and (8). Traveling wave profiles for the FKPP model (1) ($M(z)$ in blue) and the constant rate go-or-grow (8) (moving cells $U(z)$ in red, proliferating cells $V(z)$ in black). Simulations are shown at two different time steps: an earlier time $T_{1}$ (dashed curves) and a later time $T_{2}$ (solid curves). Parameters are set to $\alpha=1$, $\beta=2$, $g(0)=1$, $\mu=0$, and $d=0.0001$.
@@ 7.5   Large wave speed scaling
Introducing the traveling wave coordinate, z=xctz=x-ct, such that $U(z)=u(t,x)$ and $V(z)=v(t,x)$, the general go-or-grow model (4) is written as
@@
Example
We illustrate the large wave speed approximation in the case of the constant rates go-or-grow model (8) with $\mu=0$ and logistic growth function, g(u+v)v=(1(u+v))v{g(u+v)v=(1-(u+v))v}. In this example, system (53) becomes
$\left\{\begin{array}[]{@{}l@{}}U_{0}^{\prime}=-\alpha U_{0}+\beta V_{0},\\[11.38092pt] V_{0}^{\prime}=(1-U_{0}-V_{0})V_{0}+\alpha U_{0}-\beta V_{0}.\end{array}\right$. (54)
The classical phase plane analysis shows that the origin (0,0) is a saddle node and the equilibrium $(U_{0}^{*},V_{0}^{*})=\left(\frac{1}{\alpha/\beta+1},\frac{\alpha/\beta}{\alpha/\beta+1}\right)$} is a stable node. We calculate the nullclines of the above system in terms of $U_{0}$, obtaining
$U_{0}\mbox{-nullcline:}\qquad V_{\mathcal{N}_{1}}(U_{0})$ $=\frac{\alpha}{\beta}U_{0},$ (55a)
$V_{0}\mbox{-nullcline:}\qquad V_{\mathcal{N}_{2}}(U_{0})$ $=\frac{(1-U_{0})-\beta}{2}+\frac{\sqrt{((1-U_{0})-\beta)^{2}+4\alpha U_{0}}}{2}.$ (55b)
From these nullclines, it is possible to define the trapping region
$\mathcal{W}=\{(U_{0},V_{0})\in\mathbb{R}^{2}_{+}:0\leq U_{0}\leq U_{0}^{*},\;V_{\mathcal{N}_{1}}(U_{0})\leq V_{0}\leq V_{\mathcal{N}_{2}}(U_{0})\}\,.$ (56)
Indeed, for trajectories on the line $\{U_{0}=0,\;V_{0}>0\}$, we have $U_{0}^{\prime}>0$, thus they point inward into $\mathcal{W}$. Then, we consider the curve $\{V_{0}=V_{\mathcal{N}_{1}}(U_{0}),\;0<U_{0}<U_{0}^{*}\}$, whose trajectories have $U_{0}^{\prime}=0$ and
$V_{0}^{\prime}=\frac{\alpha}{\beta}\left[1-\left(1+\frac{\alpha}{\beta}\right)U_{0}\right]\,U_{0}>0,$
thus, the flux points upwards into $\mathcal{W}$. Finally, on the curve $\{V_{0}=V_{\mathcal{N}_{2}}(U_{0}),\;0<U_{0}<U_{0}^{*}\}$}, the trajectories have $V_{0}^{\prime}=0$ and
$U_{0}^{\prime}=-\alpha U_{0}+\beta V_{\mathcal{N}_{2}}(U_{0})>-\alpha U_{0}+\beta V_{\mathcal{N}_{1}}(U_{0})=0,$
meaning that the flux points into $\mathcal{W}$. Therefore, by Poincaré–Bendixson theorem (Theorem 3.7.1 in [96]), we have that every $\omega$-limit set of a trajectory inside $\mathcal{W}$ is either a fixed point, a periodic orbit, or a heteroclinic orbit between the two fixed points $(0,0)$ and $(U_{0}^{*},V_{0}^{*})$. Considering the Dulac function $\phi=-1/V$, we can rule out periodic orbits by applying Dulac’s criterion (Theorem 3.9.2 in [96]). Thus, we obtain the existence of a heteroclinic orbit between $(0,0)$ and $(U_{0}^{*},V_{0}^{*})$. Figure 7 illustrates the phase portrait of the system (54) and the heteroclinic orbit between the two equilibria on the left, as well as the heteroclinic orbit and the traveling wave solution of (54) in the wave coordinate $\xi$ on the right.
Figure 7: Large wave speed scaling applied to model (8). (Left) Phase portrait of system (54). Trajectories are depicted in light grey, while the blue curve represents the $V_{0}$ nullcline (55b), the red curve represents the $U_{0}$ nullcline (55a), and the black curve represents the heteroclinic connection from the $(0,0)$ equilibrium to $(U_{0}^{*},V_{0}^{*})$ (orange marker). (Right) Comparison between the traveling wave profile (solid curves) from the model (8) with the corresponding large wave speed approximation (54) (dot-dashed curves) in the coordinate $\xi$. Logistic growth is assumed in (8), and the parameters (for both plots) are set to $\alpha=1$, $\beta=0.8$, $\mu=0$, and $d=0.0001$.
@@ 8 Discussion
Focusing on specific analytical properties of go-or-grow models that can be effectively addressed, we extensively reviewed results related to the critical domain problem and the analysis of the existence and minimal speed of traveling wave solutions for the general go-or-grow model and its specialized variants. For the critical domain problem, we revisited classical results concerning the FKPP equation [9, 125] and connected them to findings related to the general go-or-grow model, which show the bifurcation between domain size and the existence of non-trivial steady-state solutions, under specific assumptions about the transition and growth functions. In our examination of traveling wave solutions, we outlined the necessary assumptions that the general go-or-grow model (4) must meet to be classified as a cooperative system. This classification allows us to apply a robust theoretical framework concerning the existence of traveling waves and invasion speed, as established in the literature [35, 128]. Similar to our analysis of the critical domain size, we establish a connection to the FKPP equation, thereby generalizing previous results to the context of the general go-or-grow model (4) introduced here.
@@ Acknowledgements
We are grateful to the Fields Institute in Toronto, who hosted us for a few weeks in Summer and Fall 2024 to work on this project. RT acknowledges support from a Canadian Graduate Scholarship of the Natural Sciences and Engineering Research Council of Canada (NSERC) and Alberta Innovates Graduate Student Scholarship. MC was supported by the National Group of Mathematical Physics (GNFM-INdAM) through the INdAM–GNFM Project (CUP E53C22001930001) “From kinetic to macroscopic models for tumor-immune system competition”, and by the European Union - NextGenerationEU and the MUR-Italian Ministry of Universities and Research through the project PRIN 2022 PNRR Mathematical modeling for a Sustainable Circular Economy in Ecosystems (project code P2022PSMT7, CUP D53D23018960001). MC has been partially supported by the State Research Agency of the Spanish Ministry of Science and FEDER-EU, project PID2022-137228OB-I00 (MICIU/AEI /10.13039/501100011033); by Modeling Nature Research Unit, Grant QUAL21-011 funded by Consejería de Universidad, Investigación e Innovación (Junta de Andalucía). TLS acknowledges support from a National Science Foundation (NSF) grant DMS-2151566. TH is supported through a discovery grant of the Natural Science and Engineering Research Council of Canada (NSERC), RGPIN-2023-04269.
@@ Appendix A Appendix: Proofs of Lemma 6 and Theorem 7
The proofs of Lemma 6 and Theorem 7 primarily require showing that the general go-or-grow model (4) with the assumptions (36b), (40a)–(40b), and (45)–(46) lives inside the general cooperative theory proposed by [77, 35, 128]. In these works, the authors are concerned the $n$-equation reaction–diffusion system
@@
[…]
@@
If the general go-or-grow model (4) satisfies (36b), (40a)–(40b), and (45)–(46), then (4) also satisfies the set of assumptions (58).
@@
Condition (58a), and the first part of condition (58c) are directly satisfied by the assumptions (36b) and (40a). In addition, during the wavefront calculation (Section 7.3), we computed the eigenvalue of the matrix $A=\rho^{2}D+{\bf F}^{\prime}(0,0)$ (42), which for $\rho=0$ reads
@@
Now, the more technical assumptions (58d) and (58e) need to be satisfied. Considering the first inequality of assumption (46), rearranging terms, and adding $\mu\,u$ to both sides yields
@@
adding $\mu\nu_{1}(\rho)$ to both sides, and multiplying by $\eta>0$, we get
@@
Therefore, if (4) satisfies (36b), (40a)–(40b), and (45)–(46), then (4) also satisfies (58). ∎
@@
  • (i) $\Phi(\mu)=\dfrac{\lambda_{A}(\mu)}{\mu}$ is decreasing as $\mu$ nears 0 and tends to infinity as $\mu\rightarrow 0$; $\Phi^{\prime}(\mu)$ changes sign at most once in $(0,\infty)$ and $\lim\limits_{\mu\rightarrow\infty}\Phi(\mu)=+\infty$;
@@
[…]
@@
Applying Lemma 10 and Theorem 11 to the general go-org-grow model (4), with assumptions (36b), (40a)–(40b), and (45)–(46), proves Lemma 6 and Theorem 7.